voluntary contributions towards a public good Shen and Antonio are considering contributing towards a creation of a water fountain each can choose whether to contribute $300 to the water fountain or to keep that $300 for a new suit.

since a water fountain is a public good both Shen and Antonio will benefit from any contributions made by the other person specifically every dollar that either one of them contribute it will bring each of them $0.90 of benefits for example if both Shen and Antonio contribute then a total of $600 would be contributed to the water fountain so then Antonio and Shen would each receive $540 of benefit from the water fountain and their combined benefits would be $1,080 this is shown in the left the first cell of the first cable

Since a new suit is a private good if Shen chooses to spend $300 on a new suit Shen would get $300 of benefit from the new suit and Antonio would receive any benefits from shins choice if Jen still spends $300 on a new suit then Antonio chooses to contribute $300 to the water fountain send would still receive the $270 of benefits from Antonio's generosity in other words it's been decides to keep $300 for a noose to dantonio decides to contribute to $300 to the public project then Shane would receive a total benefits of 300 + 270 = $570 Antonio would receive a total benefits of $270 and their combined benefits would be $840 this just shown in the lower side of the table.
complete the following

Antonio Contributes Doesnt contribute $1,080 $840 Contributes Shen Doesnt contribute of the four cells of the table, which

now consider the incentive facing Shen individually the following table looks similar to the previous one but this time it is partially completed with the individual benefits data for Shen as shown previously if both gennantonio contribute a public good Shen receive a benefit of $540 on the other hand if Antonio contributes to the water fountain and Chen does not send receives a benefit of $570.

complete the right hand column of the following table which shows the individual benefits of shen you are not required to consider the benefits of Antonio

Dont contribute Contribute $540,-. $570, -- Contribute Dont contribute Shen


fill in the blank:
if Antonio decides to contribute the water fountain send would maximize his benefit by choosing to not (contribute or to contribute)
to the water fountain
on the other hand if Antonio decides not to contribute to the water fountain Shannon would maximize its benefit by choosing( to contribute to or not contribute)
to the water fountain
these results illustrate( the common pool problem ,why markets are efficient ,the creation of positive externality, the free rider problem)
Answer #1

Benefit =300*0.9=$270

Antonio Contribute boernt. contribute . 1080 270+270+300 = 1.8401 300 + 300 840 -600 Contributes 1. Doesnt kontribute Shen I

If it helps kindly upvote

fir doubts comment below